OpenStudy (anonymous):

How do you solve for x and y-intercepts algebraically using slope-intercept and standard form?

OpenStudy (anonymous):

eg. Find the equation of the straight line that has slope m = 4 and passes through the point (–1, –6). using this equation we then solve y = mx + b (–6) = (4)(–1) + b –6 = –4 + b –2 = b Then the line equation must be y = 4x – 2

OpenStudy (anonymous):

Regardless of the form, set x = 0 to get y-intercept and set y = 0 to get x-intercept